OpenStudy (anonymous):

Not too sure about this question. A map shows the distance between the corner of Cactus Road and 1st Street and the corner of 1st Street and Merle Road as 3 inches. If the scale is 1 in. : 3.7 mi, what is the actual distance?

OpenStudy (anonymous):

Given that, Distance b/n Cactus Road Corner and 1st Street = Distance b/n 1st Street and Merle Road = 3 inches. Also, according to the scale, 1 inch = 3.7 mi So therefore, 3 inches = 3*3.7 = ?

OpenStudy (anonymous):

11.1 ?

OpenStudy (anonymous):

Yes! :) 11.1 mi is the actual distance.

OpenStudy (anonymous):

again thank you!(: